OpenStudy (anonymous):

which pair of fractions is equivalent to this pair: 3/5 and 2/7? a. 21/35 and 10/35 b. 10/35 and 7/35 c. 7/35 and 5/35 d. 24/35 and 12/35

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

3*7=21 5*7=35

OpenStudy (anonymous):

2*5=10 7*5=35